# Not-Just-Scaling Laws: Towards a Better Understanding of the Downstream Impact of Language Model Design Decisions

### v2

@@ Abstract
Improvements in language model capabilities are often attributed to increasing model size or training data, but in some cases smaller models trained on curated data or with different architectural decisions can outperform larger ones trained on more tokens. What accounts for this? To quantify the impact of these design choices, we meta-analyze 92 open-source pretrained models across a wide array of scales, including state-of-the-art open-weights models as well as less performant models and those with less conventional design decisions. We find that by incorporating features besides model size and number of training tokens, we can achieve a relative 3-28% increase in ability to predict downstream performance compared with using scale alone. Analysis of model design decisions reveal insights into data composition, such as the trade-off between language and code tasks at 15-25% code, as well as the negative impact of web data on truthfulness. Broadly, our framework lays a foundation for more systematic investigation of how model development choices shape final capabilities.  [^1] Code and data are available at https://anonymous.4open.science/r/llm-pretraining-behaviours-FE80/
[…]
@@ 1 Introduction
The research community has made progress in understanding how training decisions impact downstream performance with respect to data composition. For instance, controlled studies have demonstrated that training on code data improves performance on certain reasoning benchmarks Aryabumi et al. (2024); Petty et al. (2024), meta-features of data such as age and the use of toxicity filters affect performance on many QA tasks Longpre et al. (2024), and the balance of multilingual data affects performance on English and other languages Chang et al. (2023); Yue et al. (2025). These works uncover valuable insights, but they tend to focus on changing only a single aspect of the training recipe while keeping the rest fixed. Although rigorous, this is costly in compute and development time. We instead ask: can we leverage past findings from open language models to examine how training decisions jointly impact downstream performance?
@@
[…]
@@ 2.2 Maybe it’s Not Just Scaling?
While the majority of modern language models follow the transformer architecture, there are some details that differ. For instance, the variety (Zhang and Sennrich, 2019) and position (Xiong et al., 2020) of layer normalization, and the type of positional encoding (Su et al., 2021; Press et al., 2022) make significant differences in model performance. Previous work, such as Gu and Dao (2023), has demonstrated empirically that holding all other factors equal, models that make better architecture decisions (Touvron et al., 2023a) outperform those that make worse decisions (Vaswani et al., 2017).
@@ 3.2 Characterizing Models and Data
[…]
@@ 3.2.2 Exploring Data Composition via Generation
We propose estimating a model’s training data by generating from it with only a beginning-of-sequence token in context (or end-of-sequence if BOS is unavailable). Using temperature-based sampling with $T=1$,555this should in principle recover $P_{raw}(x_{t}\mid x_{<t})\approx P(x_{t}\mid x_{<t})$ in the limit of sampling infinitely from an LM that captures its distribution perfectly. we sample 5-10k free-generations per model and categorize them using an LM-based classifier (see Appendix E) according to the domains in Figure 2. A human annotator independently validated the classifier on 300 documents from pretraining datasets (Appendix F), showing high agreement (Cohen’s $\kappa=0.746$).
@@ 3.3 Evaluation Datasets and Metrics
Winogrande Sakaguchi et al. (2019) $\sim 44$k Accuracy
XNLI Conneau et al. (2018) $\sim 2.5$k Brier Score
@@
ARC Challenge Clark et al. (2018) $\sim 2.6$k Accuracy
Lambada Paperno et al. (2016) $\sim 10$k Accuracy
MMLU Hendrycks et al. (2020) $\sim 2.85$k Accuracy
@@ 4 Predictive Modeling
XNLI 0-shot 7.22 5.45 5.11 $\pm$ 0.11% 4.30$\pm$ 0.11% $1.37\times 10^{-2}$
ANLI 0-shot 9.48 5.95 6.18 $\pm$ 0.19% 5.86$\pm$ 0.21% $3.16\times 10^{-9}$
MathQA 0-shot 7.57 3.89 2.83 $\pm$ 0.06% 2.75$\pm$ 0.07% $1.63\times 10^{-4}$
LogiQA2 0-shot 12.62 8.77 4.74 $\pm$ 0.12% 4.60$\pm$ 0.15% $3.84\times 10^{-4}$
@@ 5.2 What Features Does Task Performance Depend On?
A little code goes a long way, but too much is harmful to NLI. The percentage of code in pretraining is a critical non-scaling feature. Higher code composition benefits Humaneval performance but negatively impacts natural language reasoning tasks including Arc Challenge, Hellaswag, Winogrande, and Lambada. As shown in Figure 5, models with over 20-25% code show gains on Humaneval but penalties on language benchmarks. A moderate 15-25% code proportion appears to balance these competing demands.
@@ 8 Conclusion and Future Work
Our current work has several limitations that can be improved in future work. First, although we document many open models, our sample size remains limited, particularly for larger ($>$50B) parameter models. This limits our ability to draw robust conclusions about scaling behaviour in large models. Additionally, the models that we have are not evenly distributed across number of parameters, data size, and data distributions, with certain size ranges and data distributions being overrepresented. There are also likely selection effects in which models are made open-weights, as well as likely time effects in popular architectural decisions or data compositions in different time periods.
@@ Appendix A List of all models
huggyllama Touvron et al. (2023a) llama-7b 7B
@@
meta-llama Touvron et al. (2023b) Llama-2-13b-hf 13B
@@ B.1 Architectural Features
  • MLP ratio - the ratio of FFN dimensionembedding $\frac{\text{FFN dimension}}{\text{embedding dimension}}$.
@@ Appendix I XGBoost Settings
To assess the impact of different data mixes on model performance, we evaluated our models on the following tasks:

### v3

# arXiv:2503.03862v3[cs.CL] 28 Feb 2026

@@ Abstract
Improvements in language model capabilities are often attributed to increasing model size or training data, but in some cases smaller models trained on curated data or with different architectural decisions can outperform larger ones trained on more tokens. What accounts for this? To quantify the impact of these design choices, we meta-analyze 92 open-source pretrained models across a wide array of scales, including state-of-the-art open-weights models as well as less performant models and those with less conventional design decisions. We find that by incorporating features besides model size and number of training tokens, we can achieve a relative 3-28% increase in ability to predict downstream performance compared with using scale alone. Analysis of model design decisions reveal insights into data composition, such as the trade-off between language and code tasks at 15-25% code, as well as the negative association of web data with truthfulness. Broadly, our framework lays a foundation for more systematic investigation of how model development choices shape final capabilities.  [^1] Code and data are available at https://github.com/nightingal3/llm-pretraining-behaviours/
@@ 1 Introduction
The research community has made progress in understanding how training decisions impact downstream performance with respect to data composition. For instance, controlled studies have demonstrated that training on code data improves performance on certain reasoning benchmarks Aryabumi et al. (2024); Petty et al. (2024), meta-features of data such as age and the use of toxicity filters affect performance on many QA tasks Longpre et al. (2024), and the balance of multilingual data affects performance on English and other languages Chang et al. (2023); Yue et al. (2025). These works uncover valuable insights, but they tend to focus on changing only a single aspect of the training recipe while keeping the rest fixed. Although rigorous, this is costly in compute and development time. We instead ask: can we leverage past findings from open language models to examine which training decisions are associated with better downstream performance?
@@
By documenting open-source models trained by the entire community and extracting insights, we provide a practical resource for model developers to learn from collective experience. We discuss this and future work in (§9)($\S\ref${sec:conclusion}).
@@ 2.2 Maybe it’s Not Just Scaling?
While the majority of modern language models follow the transformer architecture, there are some details that differ. For instance, the variety (Zhang and Sennrich, 2019) and position (Xiong et al., 2020) of layer normalization, and the type of positional encoding (Su et al., 2021; Press et al., 2022) make significant differences in model performance. Previous work, such as Gu and Dao (2023), has demonstrated empirically that holding all other factors equal, models that make better architecture decisions (Touvron et al., 2023) outperform those that make worse decisions (Vaswani et al., 2017).
@@ 3.2 Characterizing Models and Data
This reduces to typical scaling laws when $\mathcal{A}=\{\text{\# params}\}$, $\mathcal{D}=\{\text{\# tokens}\}$, and $f_{\theta}$ is a power law.
@@ 3.2.2 Exploring Data Composition via Generation
We propose estimating a model’s training data by generating from it with only a beginning-of-sequence token in context (or end-of-sequence if BOS is unavailable). Using temperature-based sampling with T=1T=1,555this should in principle recover $P_{raw}(x_{t}\mid x_{<t})\approx P(x_{t}\mid x_{<t})$ in the limit of sampling infinitely from an LM that captures its distribution perfectly. we sample 5-10k free-generations per model and categorize them using an LM-based classifier (see Appendix E) according to the domains in Figure 2. A human annotator independently validated the classifier on 300 documents from pretraining datasets (Appendix F), showing high agreement (Cohen’s $\kappa=0.746$).
@@ 3.3 Evaluation Datasets and Metrics
Winogrande Sakaguchi and others (2019) $\sim 44$k Accuracy
XNLI Conneau and others (2018) $\sim 2.5$k Brier Score
@@
ARC Challenge Clark and others (2018) $\sim 2.6$k Accuracy
Lambada Paperno and others (2016) $\sim 10$k Accuracy
MMLU Hendrycks and others (2020) $\sim 2.85$k Accuracy
@@ 4 Predictive Modeling
[…]
@@ 5.2 What Features Does Task Performance Depend On?
A little code goes a long way, but too much is harmful to NLI. The percentage of code in pretraining is a critical non-scaling feature. Higher code composition benefits Humaneval performance but is associated with lower scores on natural language reasoning tasks including Arc Challenge, Hellaswag, Winogrande, and Lambada. As shown in Figure 5, models with over 20-25% code show gains on Humaneval but penalties on language benchmarks. A moderate 15-25% code proportion appears to balance these competing demands.
@@ 6 Validating Performance Predictions with Confirmatory Experiments
7 Publication Bias Sensitivity Check
One other potential caveat to observational studies is that positive results are more likely to be reported: especially in the case of architecture, people are likely to release a new model when the model does well. There are two cases of publication bias, using architecture as an example. There are two sources of publication bias, first in design space: some architectural variants that perform particularly poorly or well may never be trained or reported; these are inherently unobservable and we cannot comment on them. We focus on the second source of bias (small-scale bias), which is that among the results that we do observe, noisier task-level estimates are more likely to look large, and thus be emphasized. We focus on this type of bias correction, which is often examined in meta-analyses.
For each default benchmark/setting in our main suite, we computed a task-level contrast (one-vs-rest for a given architectural level, e.g. RoPE vs non-RoPE positional embeddings) via precision-weighted WLS with controls for log parameters and log pretraining tokens. We then meta-regressed these task-level effects $y_{i}$ on their standard errors $SE_{i}$ via a standard method known as PET-PEESE (Stanley and Doucouliagos 2014). Across the default accuracy tasks ($k\approx 6$ per feature), the pooled, bias-corrected effects are small: RoPE $\approx\!+0.015$, RMSNorm $\approx\!+0.023$, GQA $\approx\!+0.034$ (with wider uncertainty), while other levels are near zero. Results can be found in Appendix N. These adjustments leave our qualitative story unchanged: non-scale architectural choices have, on average, modest benchmark-robust effects compared to data composition and scale.
@@ 9 Conclusion and Future Work
Our current work has several limitations that can be improved in future work. First, although we document many open models, our sample size remains limited, particularly for larger (>>50B) parameter models. This limits our ability to draw robust conclusions about scaling behaviour in large models. Additionally, the models that we have are not evenly distributed across number of parameters, data size, and data distributions, with certain size ranges and data distributions being overrepresented. There are also likely selection effects in which models are made open-weights, as well as likely time effects in popular architectural decisions or data compositions in different time periods.
@@ Appendix A List of all models
huggyllama Touvron et al. (2023) llama-7b 7B
@@
meta-llama Touvron and others (2023) Llama-2-13b-hf 13B
@@ B.1 Architectural Features
  • MLP ratio - the ratio of $\frac{\text{FFN dimension}}{\text{embedding dimension}}$.
@@ Appendix I XGBoost Settings
To assess the association of different data mixes with model performance, we evaluated our models on the following tasks:
@@
Appendix N Publication bias correction
Data from PET-PEESE on architectural variations can be found in Table 12.
Table 12: Bias-corrected architecture effects on accuracy (defaults-only; PET-PEESE). Positive values mean the feature level outperforms the rest of the pool within a task. Effects are in percentage points ($p$); CIs are 95%.
Feature Level $k$ Chosen Effect (pp) [95% CI] Egger $p$
Positional embeddings alibi 6 PEESE 2.0-2.0   [7.8-7.8, +3.9+3.9] 0.013
Positional embeddings learned 6 PET +5.4+5.4   [7.2-7.2, +17.9+17.9] 0.240
Positional embeddings ROPE 6 PEESE +1.5+1.5   [+0.9+0.9, +2.2+2.2] 0.076
LayerNorm type non-parametric 6 PEESE 3.4-3.4   [8.6-8.6, +1.8+1.8] 0.031
LayerNorm type parametric 6 PEESE 1.3-1.3   [2.7-2.7, 0.0-0.0] 0.021
LayerNorm type RMSNorm 6 PET +2.3+2.3   [+0.9+0.9, +3.8+3.8] 0.133
Attention variant full 6 PEESE +1.1+1.1   [+0.2+0.2, +2.0+2.0] 0.014
Attention variant GQA 6 PET +3.4+3.4   [+0.6+0.6, +6.3+6.3] 0.290
Attention variant local,full 6 PET +1.3+1.3   [1.8-1.8, +4.3+4.3] 0.358
Attention variant MQA 6 PEESE $0.0$   [$0.0$, $0.0$] 0.038